OpenStudy (anonymous):

Show that the number is a rational number by finding a ratio of two integers in simplest form equal to that given number. (-1/3) The number is a rational because it can be expressed as: (Help I don't get what it's asking me 2 do)

OpenStudy (zzr0ck3r):

can -1/3 be reduced?

OpenStudy (anonymous):

that's in simpelst form

OpenStudy (agent0smith):

Can you un-reduce it, then? What's another fraction equivalent to -1/3? (this may be what it's asking for)

OpenStudy (zzr0ck3r):

then the two integers that make up the number are -1 and 3

OpenStudy (zzr0ck3r):

or 1 and -3

OpenStudy (agent0smith):

Or state that it's rational, because it's in fractional form with two integers, and is simplest form.

OpenStudy (anonymous):

1/3 * -1 ?

OpenStudy (anonymous):

-1/9 * 3 ? can't really state anything much really because it's an online test

OpenStudy (zzr0ck3r):

first off read the rules, we are not here to take a test for you........

OpenStudy (zzr0ck3r):

the questions asks you to give the fraction in lowest terms, and then tell what two integers i.e. {....-3,-2,-1,0,1,2,3...} make up that fraction in lowest terms.

OpenStudy (zzr0ck3r):

the question is a bad one, because two answers.... \[-\frac{1}{3}=\frac{1}{-3}=\frac{-1}{3}\]